OpenStudy (anonymous):

Solve the following proportion. 3/y+2 = 2/y

OpenStudy (anonymous):

3y = 2(y + 2) 3y = 2y + 4 y = 4

OpenStudy (anonymous):

That's assuming that 3/y+2 is really 3/(y+2)